OpenStudy (anonymous):

I WILL MEADEL AND FAN!!!!!!!!! the length of a kitten on a scale drawing is 8cm. if the scale is 1cm:6cm, what is the actual length of the kitten? A.14cm B.42cm C.54cm D.48cm

OpenStudy (anonymous):

1 cm to 6 cm = 8 cm to x cm 1/6 = 8/x x = (6)(8) x = 48 cm So, D. (:

OpenStudy (anonymous):

thanks agian

OpenStudy (anonymous):

You're welcome.